OpenStudy (anonymous):

What is the soulotion to -(-(-)(-)(-10x)=-5

OpenStudy (lgbasallote):

-(-10x) = 10x -(10x) = -10x -(-10x) = 10x -(10x) = -10x therefore -(-(-(-(-10x ) ) ) ) = -10x so -10x = -5 10x = 5 x = 5/10 simlify

OpenStudy (anonymous):

woah woah woah slow down please can you break it down into smaller steps please

jimthompson5910 (jim_thompson5910):

We have some number x. Ten times this number is 10x The opposite of 10x is -10x

jimthompson5910 (jim_thompson5910):

The opposite of -10x is -(-10x) which is the same as 10x (basically the two opposites cancel)

jimthompson5910 (jim_thompson5910):

So each time we take the opposite of something, we either jump from positive to negative...or negative to positive

jimthompson5910 (jim_thompson5910):

So we go from -10x to 10x then we go from 10x to -10x then -10x to 10x And we're doing this five times

jimthompson5910 (jim_thompson5910):

Because we're doing it an odd number of times, the end result will be the opposite of the original...so the opposite of 10x, which is -10x So -(-(-)(-)(-10x)=-5 simplifies to -10x = -5

OpenStudy (anonymous):

werent you the person who helped me with my last equation?

jimthompson5910 (jim_thompson5910):

yes, why do you ask?

OpenStudy (anonymous):

idk the ame looked firmiliar

jimthompson5910 (jim_thompson5910):

i gotcha

OpenStudy (anonymous):

man you must pretty darn smarrt

jimthompson5910 (jim_thompson5910):

why thank you, it's from lots of practice which means that you can be just as smart someday

OpenStudy (anonymous):

r u liek in collage or somthing

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so I've basically had a lot of practice

OpenStudy (anonymous):

another question how many prime numbers are in the number 31 my lil sis wants 2 know

jimthompson5910 (jim_thompson5910):

31 is prime itself, so its only factors are 1 and 31

jimthompson5910 (jim_thompson5910):

so only one prime number factors into 31...and that's 31 itself

OpenStudy (anonymous):

srry i meant howmany prime numbers are there lass than 31

jimthompson5910 (jim_thompson5910):

oh

jimthompson5910 (jim_thompson5910):

The list of primes are... 2, 3, 5, 7, 11, 13, 17, 19, 23, 29, 31, 37, ... Of this list, only the following are less than 31 2, 3, 5, 7, 11, 13, 17, 19, 23, 29

jimthompson5910 (jim_thompson5910):

There are 10 prime numbers in that last list, so there are 10 prime numbers less than 31

OpenStudy (anonymous):

u r a life saver i probably wouldnt of understood the rest of my homerwork

jimthompson5910 (jim_thompson5910):

I'm glad I could help you out